Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 199     Entry time: 11 May 2023, 05:07
Author: Konopkova Zuzana 
Group: HED 
Subject: Shift summary 

beam parameters: 18 keV, 1 pulse on demand, 400 uJ

achieved:

  • VISAR calibrations
  • timing between X-rays, VISAR laser and DIPOLE laser
  • main shots with DIPOLE energies up to 35 J, 10 ns flat top -> shock wave compression experiments on user targets

night shift changed the phase plate to 250 um, re-aligned everything and continued with user programme succesfully

issues:

  • cryostat for the 100J is icing and needs to be defrosted which takes about 1 hour. The frequency of this happening is increasing - Over the last 24 hours it happened 3 times
  • slowness of controls - especially Karabo playground MDLs
  • slowness of DAMNIT and extraction of data tiff files
  • ADQ412 (HED_EXP_IC2/ADC/1) stops randomly acquisition on its own

X-ray beam delivery relatively stable, could make use of bit higher pulse energy. Mirror feedbacks work fine

ELOG V3.1.4-7c3fd00